OpenStudy (ggfgg):

solve the system using an augmented matrix. show all work. -x-3y=-17 2x-6y=-26

OpenStudy (holsteremission):

The augmented matrix for this system is \[\left[ \begin{array}{cc|c} -1&-3&-17\\[1ex] 2&-6&-26 \end{array} \right]\]To solve the system you need to use elimination/row reduction to end up with the augmented matrix \[\left[ \begin{array}{cc|c} 1&0&a\\[1ex] 0&1&b \end{array} \right]\]where \((a,b)\) is the solution to the system (if it exists). Do you know how to carry out row operations?